Two unknown charges attract each other. The first charge has twice greater magnitude than the second charge. If the first acting under charge is 20-12 N and the distance between them is 120 nm, what is the charge of the second charge? what is the force at 200 nm?


1
Expert's answer
2021-03-01T12:39:36-0500

"F = k\frac{q_1 \times q_2}{d^2}"

"k = 8.99 \times 10^9 \;N \;m^2/C^2" is Coulomb constants

"q_1 = 2q_2 \\

20^{-12} = 8.99 \times 10^9 \frac{2q_2^2}{(120 \times 10^{-9})^2} \\

20^{-12} = 8.99 \times 10^9 \frac{2q_2^2}{1.44 \times 10^{-14}} \\

20^{-12} = 12.48 \times 10^{23} q_2 \\

q_2 = 1.39 \times 10^{-20} \;C \\

q_1 = 2.78 \times 10^{-20} \;C"

At 200 nm:

"F = 8.99 \times 10^9\frac{2.78 \times 10^{-20} \times 1.39 \times 10^{-20}}{(200 \times 10^{-9})^2} \\

= 8.99 \times 10^9 \frac{3.86 \times 10^{-40}}{4.0 \times 10^{-14}} \\

= 8.99 \times 10^9 \times 0.965 \times 10^{-26} \\

= 8.67 \times 10^{-17} \;N"